OpenStudy (anonymous):

what is 3 over 18 written in simplest form explain how you found your answer

OpenStudy (anonymous):

3/18 is the fraction right?

OpenStudy (anonymous):

1/6

OpenStudy (anonymous):

Simplest form means that the fraction is reduce to the lowest possible numbers that it still equals the same thing. For example, \[\frac{ 5}{ 30 }\] Is about .166 This fraction isn't in simplest form because both 5, and 30 can be reduced by 5 So the simplest form would be \[\frac{ 1 }{ 6 }\] To find the simplest form, start by trying to find similar factors to both numbers (it is easiest to start at lowe numbers, like 2, and go up ill you get to the lowest number in your fraction).

OpenStudy (anonymous):

For your problem, two wouldn't work out because 3 isn't divisible by 2. So from here you would go up one, to three, and see if both numbers can be reduced (divided) by three. It turns out, both can. so the reduced fraction would be \[\frac{ 1 }{ 6 }\] which is the same to 3/18